# SOFIA/HAWC+ FAR-INFRARED POLARIMETRIC LARGE-AREA CMZ EXPLORATION (FIREPLACE) SURVEY III: FULL SURVEY DATA SET

### v2

@@ 3.1 Testing Convergence of CRUSH Reduction
The large amount of extended emission present in our observations of the CMZ requires that we run CRUSH for more than the usual 15 rounds of fitting iterations. To confirm that the CRUSH fits are converging using this larger number of rounds, we conducted a second reduction of our HAWC+ observations that used the same CRUSH parameters (e.g. using the “extended” flag and setting “-downsample=1”) but changing the number of rounds to 75 instead of 85 (“-rounds=75” instead of “-rounds=85”), matching the reduction verification conducted in FIREPLACE I. We then created difference maps of Stokes , , , and of the polarization angle () between the 85 and 75 rounds reductions to assess whether the different numbers of iterations are converging to similar solutions. The difference maps for the merged 2022 flights are shown in Figure 4.
The top left panel of Figure 4 shows the difference in Stokes between the different number of rounds. These differences are generally 1000 MJy sr, and therefore the magnitudes of these differences are 1% of the total intensity of the significant emission obtained from our observations as can be seen from the colorscale of Figure 3. The magnitudes of the differences in the Stokes (upper right panel of Figure 4) and (lower left panel of Figure 4) distributions are an order of magnitude smaller than the differences observed for Stokes , and relative to the magnitudes of the Stokes and distributions, these differences are on the 1% level in regions of high linearly polarized emission. The polarization angle differences shown in the lower right panel of Figure 4 are generally 10. Since a 10 uncertainty in polarization angle corresponds to /, the variation resulting from the different numbers of rounds is smaller than the statistical noise resulting from the significance cuts employed in this paper. We therefore conclude that changing the number of iterations from 75 to 85 rounds is not significantly changing the results obtained from the CRUSH algorithm.
@@ 3.2 Variance Within DR2 Observations
We further test the convergence of the CRUSH fits by computing the variance for the percentage fractional and parameters across the entire CMZ. The percentage fractional Stokes parameters are defined such that: . We are able to compute variance maps of our observations because our observing strategy ensures that each pixel is sampled in multiple fields. The overlap between each field is such that an image made using every third field would have contiguous, non-redundant coverage over most of the CMZ. We therefore create sub-merges using every third field and for both of the 30 orientations, resulting in 6 sub-merges of the FIREPLACE observations that cover the extent of the CMZ observed in DR2. We compute the variance for each line-of-sight in our observations by treating the measurement from the full merged data set at that sky location as the “model” value, with each of our six data subsets being the different measurements of this “model” value.
The square-root of the variance values obtained for the percentage fractional and distributions are shown in the left and middle panels of Figure 5. The right panel of Figure 5 displays the hitmap indicating how many of the six subsets of observations overlap with each pixel evaluated in this procedure. The central region of this hitmap peaks at an overlap of six, but regions towards the edge of the map exhibit less overlap because of some fields that were removed from the final set of observations due to calibration issues. In addition, there is less overlap towards the center of the hitmap as a result of an unobserved field. The number of overlapping subsets also decreases towards the edge of the distribution because of how the fields are distributed and oriented on the sky (Figure 2).
We compute the variance only for the lines-of-sight that pass the significance thresholds discussed at the end of Section 2.2. In these regions of significant total and polarized intensity, the variance is generally 1% in polarization percentage. The variance is also generally quite flat, though there are some regions where the variance is . The small percentage variances observed indicate that the Stokes , , and emission is generally consistent across the different sub-portions of the combined 2022 data set; however, some areas of the map indicate the possible presence of systematic errors as can be seen where the and variance approaches 1%.
@@ 3.3 CNM Observing Mode Comparison
To further validate the quality of our reduction we follow the procedure employed in FIREPLACE I to compare our observations with archival CNM mode observations for 3 prominent molecular clouds covered by FIREPLACE DR2: the Brick, the Circum-nuclear Disk (CND), and Sgr C. Similar to the FIREPLACE data, these CNM mode data sets are also available on the SOFIA:IRSA data archive. The CNM mode data set for the Brick, consisting of 5 files, was obtained on flight 394 on May 12, 2017 by principle investigator T. Pillai (proposal ID: 0206). The CNM mode data for the CND and Sgr C regions were obtained on flight 397 on May 18, 2017 by principle investigator D. Chuss (proposal ID: 0018), consisting of 4 files for the CND and 3 files for Sgr C.
@@
Figure 6 (left column) shows the inferred polarization pseudovectors of the CNM mode and OTFMAP mode data in (top) the Brick, (middle) the CND, and (bottom) Sgr C. The pseudovectors in these maps have been shifted slightly for ease of comparison. A polarimetric signal-to-noise cut for each data set, , and a total intensity signal-to-noise cut for the OTFMAP mode data set, , were applied, which are the same significance cuts employed in Section 2.2. The polarization angles of the CNM mode and OTFMAP mode observations generally agree for intensities greater than the red dashed vertical line in the right hand panels of Figure 6. Less agreement is seen around low intensity regions for all three clouds where the reference beam contamination for the CNM mode observations is potentially more significant (i.e. the spatial differences between the two observing methods is most impactful).
@@
The right column of Figure 6 quantifies the comparison through calculation of the Pearson correlation between the normalized Stokes and parameters of the CNM mode and OTFMAP mode data as a function of a total intensity lower threshold. The Pearson- value is calculated only between points where the total intensity exceeds the intensity threshold, as determined from the OTFMAP mode total intensity map. The maximum intensity value was chosen to ensure that a minimum of 100 pixels (approximately 6 independent beams) are included in each correlation.
@@
For intensities less than the red dashed vertical line in the right panels of Figure 6 the correlations are observed to drop. This drop is likely due to differences in spatial filtering between the OTFMAP and CNM mode observations, since spatial filtering is more significant in lower total intensity regions. At higher total intensities there are discrepancies between the Stokes parameters where one of the normalized Stokes parameters has a higher Pearson correlation than the other. For example, the correlations derived for the CND reveal a systematically higher correlation for normalized fractional Stokes than for normalized fractional Stokes . This discrepancy is caused by different signal-to-noise levels for the different normalized parameters where in the case for the CND there is generally much higher signal in than in .
In general, for all three clouds there is an increase in the Pearson correlation with increasing intensity. This trend is expected, since at higher total intensities contamination in the reference beam is less significant. The trend of increasing correlation with total intensity supports the validity of the data reduction pipeline employed for our FIREPLACE observations.
@@ 3.4 Comparison With Herschel 250 µm
As an additional check of the consistency of our new FIREPLACE observations, we plot the FIREPLACE DR2 Stokes (214 µm intensity) against the Herschel 250 µm intensity (Molinari et al., 2011) in Figure 7. The Pearson value is found to be 0.91, showing good correlation between the two data sets.
There are instances where the 214 µm intensity from FIREPLACE is 0 as can be seen in Figure 7. The negative emission indicates that the SOFIA/HAWC+ observations are insensitive to large-scale emission associated with very extended structures.
Though there is variation in dust temperature throughout the CMZ, we note that the slope of 0.91 found is consistent with dust having a temperature of 12 K with a spectral index of . This temperature is low for the molecular population in the GC, which exhibit an average temperature of 20 K (Molinari et al., 2011). This estimate of the slope is biased downward by the fact that the SOFIA/HAWC+ Stokes map is less sensitive to extended emission than the Herschel 250 µm map as described previously when discussing the negative 214 µm intensities observed in Figure 7.
Fitting only the positive values in Figure 7 would yield a larger slope for the fit to the correlation between FIREPLACE and Herschel. A higher slope would be consistent with dust possessing a higher temperature, and would therefore be more aligned with the average molecular temperature observed in the CMZ.
@@ 4 The Polarimetric CMZ at 214 µm
Having verified our DR2 reduction we combine these DR2 data with the DR1 observations presented in FIREPLACE I for a full map of the CMZ. This combination is straightforward since the DR1 and DR2 fields were reduced using identical CRUSH parameters. We therefore co-added all of the individual DR1 and DR2 fields using the DRP merge algorithm. The combined fields consist of the DR2 fields presented previously in Table A.2.1 and the set of DR1 fields shown in Table A.2.2.
@@
The upper panel of Figure 8 reveals the polarization pseudovectors (yellow dashes) obtained from our combined FIREPLACE DR1 and DR2 observations that cover the entire CMZ. Polarization orientations are only displayed for lines of sight that pass the selection criteria described in Section 2.2. Even with these significance cuts we obtain 65,000 Nyquist-sampled pseudovectors from our full 214 µm map of the CMZ.
@@
Following the approach of FIREPLACE I, for the remainder of this paper we assume that the dust grain alignment mechanism responsible for the polarized emission is such that the angular momentum vector of the grains is predominantly parallel to the magnetic field direction and that the grains are spinning preferentially about their axis of greatest moment of inertia. This implies that the inferred direction of the magnetic field in the plane of the sky is perpendicular to the emergent polarization (or electric field) direction. This is consistent with -RAT alignment theory (Andersson et al., 2015; Lazarian & Hoang, 2007). We note that ultimately this interpretation of polarization will rely on a more complete theoretical understanding of the physics of grain alignment than currently exists. The lower panel of Figure 8 shows the derived magnetic field pseudovectors obtained by rotating the polarization angle orientations shown in the upper panel of Figure 8 by 90 under the assumption of -RAT alignment.
@@
We display the Line Integral Contour (LIC) representation of these magnetic field pseudovectors in Figure 9, made using a kernel size of 1′. The LIC is a method of illustrating the polarization pseudovectors as streamlines (Cabral & Leedom, 1993). To do so, we filter the input distribution of polarization pseudovectors along local streamlines to produce an output image that represents the ‘flow’ direction of the polarization orientations. This visualization technique leads to a representation of the polarization that emphasizes the field geometry on the scales of molecular clouds in the CMZ.
@@
The magnetic field orientation obtained from FIREPLACE is seen to generally trace the morphology of individual molecular clouds, as evident in the inset panels of Figure 9. The fact that the magnetic field orientation is impacted by the morphology of the CMZ molecular clouds indicates that the field derived in FIREPLACE is likely local to the CMZ rather than some other magnetic field component encountered along the line-of-sight.
@@
A comparison of the FIREPLACE sample to previous infrared polarimetric data sets of individual molecular clouds within the CMZ made using the Kuiper Airborne Observatory (KAO), Caltech Submillimeter Observatory (CSO), and James Clerk Maxwell Telescope (JCMT) is shown in Figure 10. THE KAO and CSO pseudovectors are scaled like the FIREPLACE pseudovectors: proportional to fractional polarization. The JCMT pseudovectors, conversely, are not scaled to match their presentation in Lu et al. (2023). It is clear from Figure 10 that the FIREPLACE survey provides extensive polarimetric coverage of the CMZ, yielding a large sample of pseudovectors that will be used for statistical analysis. Figure 10 clearly reveals the sensitivity improvement the FIREPLACE observations provide in probing the magnetic field local to the CMZ molecular cloud population. In particular, the FIREPLACE observations yield sensitivity to fainter dust emission regions further from the Galactic plane and between the prominent molecular clouds in the CMZ. This increased sensitivity is a result of our observing strategy utilizing OTFMAP mode polarimetry and overlapping fields. The FIREPLACE magnetic field directions are in general agreement with those of previous infrared polarimetric observations of prominent CMZ molecular clouds, as can be seen in Figure 10; however, there are examples where the field derived from FIREPLACE is systematically different than previous observations. Comparisons between the pseudovectors derived from the FIREPLACE observations and previous individual cloud observations are discussed in more detail in Section 5.
@@ 4.1 The CMZ Magnetic Field on Multiple Spatial Scales
We first compare the magnetic field orientations obtained from FIREPLACE at 19.6″ resolution with the magnetic fields derived from ACTPol at 1′ and PILOT at 2′ resolution (Guan et al., 2021; Mangilli et al., 2019, respectively). This comparison is shown in Figure 11 where the magnetic fields are superposed on the 250 µm total intensity () CMZ emission from Herschel (Molinari et al., 2011). In general the magnetic fields from PILOT and ACTPol are smooth and parallel to one another, indicating that they are tracing a similar magnetic field component. The magnetic field traced by FIREPLACE is more spatially variant than either the PILOT or ACTPol distributions. The FIREPLACE magnetic field tends to trace the total intensity morphology of prominent CMZ molecular clouds as seen in high emissivity regions in Figure 11 and previously in Figure 9. In lower emissivity regions the magnetic field does not exhibit an apparent preferential orientation, ranging from being parallel to the Galactic plane to perpendicular to the Galactic plane.
Because of the sensitivity of ACTPol and PILOT to large-scale polarization relative to FIREPLACE, the magnetic fields derived from these surveys likely includes a magnetic field contribution originating from emission external to the CMZ. In other words, the increased spatial filtering and higher resolution of FIREPLACE relative to PILOT and ACTPol likely results in measurements of polarization local to the dust structures in the CMZ, whereas the PILOT and ACTPol measurements include contributions from a larger part of the central kpc of the Galaxy. Careful modeling of potential line-of-sight magnetic field components for these data sets will enable an analysis of the magnetic field configuration along the line-of-sight. For example, comparison of IR emission from stars on the near and far side of the CMZ indicates that this field component is connected to material that is within the central 12 kpc of the Galaxy (Nishiyama et al., 2009).

### v3

@@ 3.1 Testing Convergence of CRUSH Reduction
The large amount of extended emission present in our observations of the CMZ requires that we run CRUSH for more than the default 15 rounds of fitting iterations. To confirm that the CRUSH fits are converging using this larger number of rounds, we performed a second reduction of our HAWC+ observations that used the same CRUSH parameters (e.g. using the “-extended” flag and setting “-downsample=1”) but changing the number of rounds to 75 instead of 85 (“-rounds=75” instead of “-rounds=85”), matching the reduction verification conducted in FIREPLACE I. We then created difference maps of Stokes , , , and of the polarization angle () between the 85 and 75 rounds reductions to assess whether the different numbers of iterations are converging to similar solutions. The difference maps for the merged 2022 flights are shown in Figure 4.
The top left panel of Figure 4 shows the difference in Stokes between the different number of rounds as a percentage difference. These differences are generally 1% in the region satisfying the significance cuts described in Section 2.2. A percentage difference of 1% corresponds to a magnitude difference of 1000 MJy sr-1 in Stokes . The magnitudes of the differences in the Stokes (upper right panel of Figure 4) and (lower left panel of Figure 4) distributions are on the order of 100 MJy sr-1, corresponding to 1% the and intensities observed in the significance region. Because Stokes and are angle-based parameters, we express the differences in and as absolute differences rather than percentage differences in Figure 4. The polarization angle differences shown in the lower right panel of Figure 4 are generally 10∘. Since a 10∘ uncertainty in polarization angle corresponds to /, the variation resulting from the different numbers of rounds is smaller than the statistical noise resulting from the significance cuts employed in this paper. We therefore conclude that changing the number of iterations from 75 to 85 rounds is not significantly changing the results obtained from the CRUSH algorithm.
@@ 3.2 Variance Within DR2 Observations
We further test the convergence of the CRUSH reduction by computing the square-root of the variance for the fractional and parameters across the entire CMZ. The fractional Stokes parameters (in percentage) are defined such that: . We are able to compute variance maps of our observations because our observing strategy ensures that each pixel is sampled in multiple fields. The overlap between each field is such that an image made using every third field would have contiguous, non-redundant coverage over most of the CMZ. We therefore create sub-merges using every third field and for both of the 30∘ scan orientations, resulting in 6 sub-merges of the FIREPLACE observations that cover the extent of the CMZ observed in DR2. We compute the variance for each line-of-sight in our observations by treating the measurement from the full merged data set at that sky location as the “model” value, with each of our six data subsets being the different measurements of this “model” value.
The square-root of the variance values obtained for the fractional and distributions are shown in the left and middle panels of Figure 5. The right panel of Figure 5 displays the hitmap indicating how many of the six subsets of observations overlap with each pixel evaluated in this procedure. The central region of this hitmap peaks at an overlap of six. The number of overlapping subsets decreases towards the edge of the map because of how the fields are distributed and oriented on the sky (Figure 2).
We compute the variance only for the lines-of-sight that pass the significance thresholds discussed at the end of Section 2.2. In these regions of significant total and polarized intensity, the square root of the variance is generally 1%. The variance is also generally quite flat, though there are some regions where the variance is . The small percentage variances observed indicate that the Stokes , , and emission is generally consistent across the different sub-portions of the combined 2022 data set. Furthermore, these variances are of the order of the reported instrumental polarization of the HAWC+ instrument (2%), indicating the impact of polarization leakage in high emission regions may be negligible. Some areas of the map indicate the possible presence of systematic errors, however, as can be seen where the and square root variance approaches 1%.
@@ 3.3 CNM Observing Mode Comparison
To further validate the quality of our reduction, we follow the procedure employed in FIREPLACE I to compare our observations with archival CNM mode observations for 3 prominent molecular clouds covered by FIREPLACE DR2: the Brick, the Circum-nuclear Disk (CND), and Sgr C. These CNM mode data sets are available on the SOFIA:IRSA data archive. The CNM mode data set for the Brick, consisting of 5 files, was obtained on flight 394 on May 12, 2017 by Principal Investigator T. Pillai (proposal ID: 05_0206). The CNM mode data for the CND and Sgr C regions were obtained on flight 397 on May 18, 2017 by Principal Investigator D. Chuss (proposal ID: 05_0018), consisting of 4 files for the CND and 3 files for Sgr C.
@@
Figure 6 (left column) shows the inferred polarization pseudovectors of the CNM mode (blue pseudovectors) and OTFMAP mode (red pseudovectors) data in the Brick (top), the CND (middle), and Sgr C (bottom). The pseudovectors in these maps have been shifted by 1 pixel (5″) from one another for ease of comparison. We applied the same significance cuts to the CNM mode observations that were applied to our OTFMAP mode observations discussed in Section 2.2 of , 50%, and to preserve consistency between the two observing modes. The polarization angles of the CNM mode and OTFMAP mode observations generally agree for intensities greater than the red dashed vertical line in the middle and right columns of Figure 6. Less agreement is seen around low intensity regions for all three clouds where the reference beam contamination for the CNM mode observations is potentially more significant (i.e. the spatial filtering differences between the two observing methods is most impactful).
@@
The middle column of Figure 6 quantifies the comparison through calculation of the Pearson correlation between the normalized Stokes and parameters of the CNM mode and OTFMAP mode data as a function of a total intensity threshold. The Pearson- value is calculated only between points where the total intensity exceeds the intensity threshold, as determined from the OTFMAP mode total intensity map. The maximum intensity value was chosen to ensure that a minimum of 100 pixels (approximately 6 independent beams) are included in each correlation.
@@
For intensities less than the red dashed vertical line in the right panels of Figure 6 the correlations are observed to drop. This drop is likely due to differences in spatial filtering between the OTFMAP and CNM mode observations, since spatial filtering is more significant in lower total intensity regions. At higher total intensities there are discrepancies between the Stokes parameters where one of the normalized Stokes parameters has a higher Pearson correlation than the other. For example, the correlations derived for the CND reveal a systematically higher correlation for normalized fractional Stokes than for normalized fractional Stokes . This discrepancy is caused by different signal-to-noise levels for the different normalized parameters where in the case for the CND there is generally much higher signal in than in . In general, for all three clouds there is an increase in the Pearson correlation with increasing intensity. This trend is expected, since at higher total intensities contamination in the reference beam is less significant.
The right column of Figure 6 displays the slope of the and parameters vs intensity threshold. For all three clouds we observe the slope approach a value of 1. The slope of 1 indicates we are finding similar polarization fractions between the CNM and OTFMAP mode observations. Such a result indicates that the differences we observe between the CNM and OFTMAP mode observations are not a result of a systematic offset between the different observations, but rather differences in spatial filtering or potential reference beam contamination in the CNM mode observations. The trend of increasing correlation with total intensity and the and slopes of 1 support the validity of the data reduction pipeline employed for our FIREPLACE observations.
@@ 3.4 Comparison With Herschel 250 µm
As an additional check of the consistency of our new FIREPLACE observations, we plot the FIREPLACE DR2 Stokes (214 µm intensity) against the Herschel 250 µm intensity (Molinari et al., 2011) in Figure 7. The Pearson value is found to be 0.89, showing good correlation between the two data sets.
There are instances where the 214 µm intensity from FIREPLACE is 0 as can be seen in Figure 7. The negative emission indicates that the SOFIA/HAWC+ observations are insensitive to large-scale emission associated with very extended structures. Though there is variation in dust temperature throughout the CMZ, we note that the slope of 1.14 found is consistent with dust having a temperature of 17.5 K with a spectral index of . This temperature is in good agreement for the dust in the GC, which exhibits an average temperature of 20 K (Molinari et al., 2011).
The SOFIA/HAWC+ Stokes map is possibly less sensitive to extended emission than the Herschel 250 µm. This loss in sensitivity could explain the negative 214 µm intensities observed in Figure 7. Fitting only the positive values in Figure 7 yields essentially the same values for Pearson r and slope, indicating that the negative values are not significantly impacting our correlation with the Herschel 250 µm observations.
@@ 4 The Polarimetric CMZ at 214 µm
Having verified our DR2 reduction, we combine these DR2 data with the DR1 observations presented in FIREPLACE I to obtain a full map of the CMZ. This combination is straightforward since the DR1 and DR2 fields were reduced using identical CRUSH parameters. We therefore co-added all of the individual DR1 and DR2 fields using the DRP merge algorithm. The combined fields consist of the DR2 fields presented previously in Table A.1 and the set of DR1 fields shown in Table A.2.
@@
The upper panel of Figure 8 reveals the polarization pseudovectors (yellow pseudovectors) obtained from our combined FIREPLACE DR1 and DR2 observations that cover the entire CMZ. Polarization orientations are only displayed for lines-of-sight that pass the selection criteria described in Section 2.2. Even with these significance cuts we obtain 64,000 Nyquist-sampled pseudovectors from our full 214 µm map of the CMZ, a significant increase from the 24,000 pseudovectors obtained in DR1 (FIREPLACE I).
@@
Following the approach of FIREPLACE I we assume that the dust grain alignment mechanism responsible for the polarized emission results from the grains preferentially aligning perpendicular to the magnetic field orientation. The dust grains spin preferentially about their axis of greatest moment of inertia. This scenario implies that the inferred direction of the magnetic field in the plane of the sky is perpendicular to the emergent polarization (or electric field) direction. This is consistent with -RAT alignment theory (Lazarian & Hoang, 2007; Andersson et al., 2015). We note that ultimately this interpretation of polarization will rely on a more complete theoretical understanding of the physics of grain alignment than currently exists.
The lower panel of Figure 8 shows the derived magnetic field pseudovectors obtained by rotating the polarization angle pseudovectors shown in the upper panel of Figure 8 by 90∘ under the assumption of -RAT alignment. These magnetic field pseudovectors are shown overlayed on the FIREPLACE emission.
@@
We display the Line Integral Contour (LIC) representation of the FIREPLACE magnetic field pseudovectors in Figure 9, made using a kernel size of 1′. The LIC is a method of illustrating the magnetic field pseudovectors as streamlines (Cabral & Leedom, 1993). To do so, we filter the input distribution of magnetic field pseudovectors along local streamlines to produce an output image that represents the ‘flow’ direction of the magnetic field orientations. This visualization technique leads to a representation of the magnetic field that emphasizes the field geometry on the scales of molecular clouds in the CMZ.
@@
The magnetic field orientations obtained from FIREPLACE are seen to generally trace the morphology of individual molecular clouds, as evident in the inset panels of Figure 9. The fact that the magnetic field is impacted by the morphology of the CMZ molecular clouds indicates that the field derived from FIREPLACE is likely local to the CMZ rather than some other magnetic field component encountered along the line-of-sight. This conclusion is consistent with the idea that polarized thermal emission is column density weighted within the volume of the beam (e.g., Reissl et al., 2020).
@@
A comparison of the FIREPLACE pseudovectors to previous infrared polarimetric data sets of individual molecular clouds within the CMZ made using the Kuiper Airborne Observatory (KAO), Caltech Submillimeter Observatory (CSO), and James Clerk Maxwell Telescope (JCMT) is shown in Figure 10. The KAO and CSO pseudovectors are scaled in a manner similar to the FIREPLACE pseudovectors: proportional to fractional polarization. The JCMT pseudovectors, conversely, are not scaled, which matches their presentation in Lu et al. (2024). It is clear from Figure 10 that the FIREPLACE survey provides extensive polarimetric coverage of the CMZ, yielding a large sample of pseudovectors that will be used for statistical analysis. Figure 10 clearly reveals the ability of the FIREPLACE observations to measure the magnetic field in lower intensity regimes. In particular, the FIREPLACE observations recover pseudovectors corresponding to fainter dust emission regions further from the Galactic plane and between the prominent molecular clouds in the CMZ compared to these other observations. This increased sensitivity is a result of our observing strategy utilizing OTFMAP mode polarimetry and overlapping fields. The FIREPLACE magnetic field orientations are in general agreement with those of previous infrared polarimetric observations of prominent CMZ molecular clouds, as can be seen in Figure 10; however, there are examples where the field derived from FIREPLACE is systematically different than previous observations. Comparisons between the pseudovectors derived from the FIREPLACE observations and previous individual cloud observations are discussed in more detail in Section 5.
@@ 4.1 The CMZ Magnetic Field on Multiple Spatial Scales
We first compare the magnetic field orientations obtained from FIREPLACE (19.6″) with the magnetic fields derived from ACTPol smoothed to 2′ (Guan et al., 2021) and PILOT (also 2 ′, Mangilli et al., 2019). This comparison is shown in Figure 11 where the magnetic fields are superposed on the 250 µm total intensity () CMZ emission from Herschel (Molinari et al., 2011). In this figure, the pseudovectors have not been scaled by fractional polarization to make it easier to compare the orientations observed in the different observations. In general, the magnetic fields from PILOT and ACTPol have little variation between neighboring pseudovectors and are parallel to one another, indicating that they are tracing a similar large-scale, spatially uniform, magnetic field component. The magnetic field traced by FIREPLACE shows more spatial variation than either the PILOT or ACTPol distributions. The FIREPLACE magnetic field tends to trace the total intensity morphology of prominent CMZ molecular clouds as seen in high emissivity regions in Figure 11 and previously in Figure 9. In lower emissivity regions the magnetic field does not exhibit an apparent preferential orientation, ranging from being parallel to the Galactic plane to perpendicular to the Galactic plane.
Because of the higher sensitivity of ACTPol and PILOT to large-scale polarization patterns relative to FIREPLACE, the magnetic fields derived from these surveys likely include a magnetic field contribution originating from emission external to the CMZ. In other words, the increased spatial filtering and higher resolution of FIREPLACE, relative to PILOT and ACTPol, likely results in measurements of polarization local to the dust structures in the CMZ, whereas the PILOT and ACTPol measurements include contributions from a larger part of the central kpc of the Galaxy. Different dust temperatures encountered along the line-of-sight for the surveys is an alternative explanation for the different magnetic field orientations. However, the similarity of the field orientation observed between PILOT at 240 µm and ACTPol at 1.36 mm potentially indicates that different dust temperatures is not a significant factor in the different magnetic field orientations. Careful modeling of potential line-of-sight magnetic field components for these data sets will enable an analysis of the magnetic field configuration along the line-of-sight. For example, comparison of IR emission from stars on the near and far side of the CMZ indicates that this large-scale field component is connected to material that is within the central 12 kpc of the Galaxy (Nishiyama et al., 2009).